# Attention is All They Need: Exploring the Media Archaeology of the Computer Vision Research Paper

### v2

@@ 4.1.2. Beyond the paper: arXiv, social media and videos
Why would many authors begin to engage in these teaser figure and title practices over the course of the 2010s? We find that it has resulted from a change in medium through which research reaches its audience, from conference and journal proceedings onto social media and preprint websites. P4 describes how his students have strategies to get their papers noticed on the preprint site arXiv: “if they take your submission on Thursday and it goes up on the arXiv most recent publications list it’s up for all of Friday, Saturday, and Sunday…so it gets more days of exposure…It improves your odds that somebody will notice your thing.” P8 describes the pressure on her students to promote their work: “social media, like promoting one’s research has become such a big thing and I think students…are realizing, oh it’s not enough for me to, y’know, come up with a paper, post it on arXiv, get it accepted. I also need to tweet about it. It’s frankly quite exhausting…[before] the only way you promote your paper is it shows up at the conference and hopefully, y’know, some famous computer vision researchers will come up and look.” She talks about how her advisor would bring his friends over to her poster, and that was all the advertising she needed to get noticed, but that strategy no longer works as well in the modern crowded field of computer vision.
@@
That attention translates into citations, as well as other benefits. P11, who works in an industry lab that hires recent PhDs, explains why his group publishes at all: “exposure is actually really important. If you want to attract really top level talent, then having zero published papers is really going to work against you, right? Particularly if you’re looking at people who are in positions that are for doing some sort of active innovation.” For this group, the publication is primarily a way to gather attention, they derive value from the attention paid to their work through hiring. Second, attention helps researchers to identify the value of their own work. P7 talks at length about his desire for impact. He switched fields from computational complexity to computer vision because there were more active researchers, and he could not tell if his work was having an impact in a small field. P9 tells the story of how attention from a senior scholar improved his relationship with his advisor: “that email from [scholar], that’s the beginning of our of my great relationship with my advisor was that one little email, because within five minutes [advisor] was up in the office, ‘Oh, you got an email from [scholar]? Very good. Very good. How about a coffee?’ ” Notice that the career benefits resulted, not from the attention itself, but from the record of that attention in the form of an email from a senior scholar.
While the relationship between attention and prestige is not new, the unprecedented growth of computer vision over the past decade places pressure on the attention of senior researchers. There have been ongoing discussions in the computer vision community on how to react to this growth. Some of the proposed ideas are recorded in the passed motions of the annual meeting of the IEEE pattern analysis and machine intelligence technical committee (PAMI-TC), the governing body of the computer vision research community. In 2021-2022, motions changed the rules regarding double submission to reduce burden on reviewers, introduced punitive measures for missing or poor quality peer reviews from authors and banned discussion of submitted papers on social media to avoid influencing peer reviewers, the last of which was repealed in 2023 (pam, 2024; Black, 2021). The rationale statement for the social media ban specifically highlights attention: “Groups with large followings and the resources to mount visible social media promotions received significant attention for work that is under review. Reviewers are exposed to this work and the attention it receives can bias their judgment—if so many people on social media are excited, mustn’t it be good?” (Black, 2021) In addition to highlighting the valorizing role of attention, this statement showcases the view of peer review as an unbiased process for evaluating the quality of research.
@@ 4.2. Measuring the Contribution
But all of this was not the case a few decades ago. The vast majority of computer vision papers from the 1980s and 1990s rely on mathematical arguments based on pinhole camera geometry and do not contain any quantitative results. Empirical evaluation, if included at all, was primarily qualitative, in the form of figures showing sample results. As P8 explains, she had a combination of quantitative and qualitative evaluation in her paper from 2003, which was unusual: “quantitative evaluation, you know, back in 2003 was still kind of in its infancy…I’m not sure that this [2003] paper has basically any comparison to competing methods which probably would be required today.” P3 explains that he was primarily concerned in 1990 with showing test examples to demonstrate his algorithm’s effective handling of edge cases. P9 explains that in 1999, showing example output of his system was sufficient: “instead of [Amazon] Mechanical Turk you just have the reviewers just eyeball the images.” In the satirical 2010 “Paper Gestalt” (Von Bearnensquash, 2010) paper, which attempts to use computer vision methods to distinguish between good and bad papers, large confusing tables were identified as a key feature of bad papers, not an essential feature of good ones.
@@
The publication of Krizhevsky, Sutskever and Hinton’s “ImageNet Classification with Deep Convolutional Neural Networks” in 2012 (Krizhevsky et al., 2012) marks a turning point for empirical evaluation. This paper is historically significant for setting off the deep learning revolution, and its design and writing served as a foundation for the thousands of deep learning-based computer vision papers that followed. The paper’s central argument is that several “new and unusual features” lead deep convolutional neural networks to significantly outperform other methods. These features include rectified linear units (ReLU), GPU-based training, and regularization techniques like data augmentation and dropout. Crucially for our story, however, this argument is made by way of a table, shown in Figure 5 (c), with the best performance in bold. Neural network papers are obligated to use empirical evaluation, as there are insufficient theoretical guarantees for these models and they are difficult to evaluate otherwise. Over the following years, many papers followed Krizhevsky’s lead, showing that deep convolutional neural networks outperform existing methods on other central problems like object detection and semantic segmentation. We can see a corresponding increase in both the average number of tables per paper, as well as the fraction of papers containing at least one table in Figure 4. While the prevalence of figures has remained relatively constant, tables have become significantly more common. While only 75% of CVPR 2013 papers had a table, 95% of CVPR 2021 papers did, and the average number of tables per paper has doubled from 2 to 4. While not all of these tables are the kind of results table we’re discussing, the change is striking.
@@
More broadly, in computer vision there is widespread assumption that research inherently involves competition between technical methods. Several of our participants described academic publication using free market metaphors. P2 explains his research output: “publication was really fast during that time because there was not a lot of competition.” P3 explained that CVPR has become inaccessible to students because “supply and demand” have raised the standards for publication. P4 described vision as “very industrial” and gave examples of techniques that students use to optimize their arXiv submissions to reach as many eyes as possible.
While these tables are buried deep within the paper, it may seem like their effect on attention is minimal. However, a competitive benchmark result attracts attention on its own. Once a researcher has the lead, other papers must compare against their benchmark and cite them. P6 compares benchmarks to arcade game leaderboards: “If it’s an established benchmark, there’s like there’s somebody who has the lead right? Like you would have on like a classic video game arcade machine, right? It’s like I have to just get my initials at the top right? That’s exactly what they’re doing, and that’s frustrating.” P5 and P8 advise their students to avoid working on research which forces them into competition with large companies. P5: “I tell them, don’t work on problems that, you know, a lot of people are working on right now, you can’t possibly compete with Facebook, Google, Amazon because you’re not as computing heavy.” While students could still develop innovative ideas for those problems, they likely would not beat the benchmarks that large companies have set without leveraging similar computing power and would not attract attention.
In summary, the development of the results table shows the transition of computer vision from an applied mathematical discipline to a quantitative-empirical one. As vision started relying on empirical evaluation, it adopted a style of results table which was used for runtime benchmarking in computer graphics, and as evaluation benchmarks became established for vision problems, these tables grew in size and importance. Today, competition on benchmarks is the organizing principle of the discipline, new methods must demonstrate that they are empirically more effective than existing methods to be accepted, and the results table is an essential part of the paper. This element showcases several of the same themes as teaser images: influence from computer graphics, and commodification, in the form of measurable improvements over prior work. It also echoes patterns observed in HCI more broadly regarding quantification: once a phenomenon, in this case the quality of a method, has been measured, that measurement creates and constrains possibilities for action (Pine and Liboiron, 2015).
@@ 4.3.1. The PDF and digital proceedings
The final element we explore is the Portable Document Format (PDF) itself, and the way that authors engage with it as a medium. In this section, to avoid confusion, we use “research paper” to refer to printed papers and “research PDF” to refer to the digital version. PDF was a derivative of Adobe’s already-successful Postscript programming language for specifying documents, and a proprietary file format until 2008. In an article from 1998, Kasdorf compares the PDF format to SGML (Standard Generalized Markup Language), and argues that both formats should be used on the web: markup languages for screen-based content and page-based languages for printed content (Kasdorf, 1998).
@@
One such choice is the high resolution figure, which exploits the scalable nature of vector graphics and the zoom feature of PDF viewers to put more data in a research PDF without exceeding the page limit. Several examples of this trend from CVPR 2021 are shown in Figure 6. While these figures are not overly complex or poorly designed, they are too small to be readable on a 300-dpi printed page.
@@
Several of our participants mentioned these high resolution figures. P6 and P10 are frustrated that they cannot print some PDFs because the figures are too small. P10 ascribes these figures to the increasing complexity of research: “the tendency of grouping more small figures into a kind of collection… I wonder, is this because of the limited page count that people try to avoid the whitespace around the figures and then they put everything in one big figure because then you can save on the whitespace? Or is it simply because really…the techniques that we’re describing nowadays are so, so much more complicated? So you need them to, you know, to, to put the architecture of the whole thing in detail there. Otherwise nobody understands it.” P7, on the other hand, ascribes it to digital submissions: “In ’94, you turned in your paper by actually having a piece of large paper that was probably 24” wide by 30” or something… you had columns of text from LaTeX and then your figures had to be glued in place in what’s called ‘screen,’ right? So you actually have to make sure that they didn’t copy badly.” When papers were submitted by mail, the medium of the “camera-ready” draft was a significant limitation on the kinds of visual content that could be included, but now the only limitation is on the number of pages, so authors use smaller figures to include more content. The small size of these figures, especially in comparison with the large front page teaser image, creates a visual hierarchy, highlighting the differences in purpose between these figures. The teaser attracts attention, the dense figures provide details for those who zoom in, and the impression of details for those who do not.
@@
The digital nature of contemporary research PDFs was a sensitive topic in our interviews. Half of our participants (P1, P2, P3, P4, P5, and P8) expressed attachment to and nostalgia for their printed papers, and were saddened when asked when they stopped reading on paper. For P1, the time spent at the library made the papers physically significant: “[the papers] were all me standing at a Xerox machine with a journal…I didn’t find it as annoying. I found that it gave me something like a kind of a physical connection to the paper. This was sort of like, it’s mine. I’ve watched it go by, you know.” P2 remembers that the SIGGRAPH proceedings “were like flipped to the point where like it’s the threads were like bare.” He’s kept all of the papers he printed during his master’s degree and revisits them from time to time: “I still go back, you know, look at like oh this crap that I thought about these things, right?” Finding and photocopying papers made them precious; he thinks there is “a kind of a correlation between that sense of, like scarcity of it, like how precious the paper was, because it’s so hard to find versus the amount of care you give it.” P4 fondly recalls his advisor taking papers out of a filing cabinet and photocopying them for him: “the amazing thing about your advisor is you have this filing cabinet of stuff that’s already pre-copied, you know, every paper he thought was interesting.”
@@
These sorts of relationships to printed papers are interesting from a design perspective. They echo the findings of Odom et al. that we preserve designed objects which have functional, symbolic and material-aesthetic value (Odom et al., 2009). When a researcher prints out and writes on a paper, the paper gains symbolic significance, a “physical connection” which endures over time. That makes printed pages less disposable than digital ones; there is a more tangible opportunity cost to creating them. From this perspective, digitization is an essential component of commodification, as it separates the research from its paper and thus its role as symbolic object. While reading digitally is significantly more environmentally sustainable than printing, shipping and photocopying paper (Aydemir and Özsoy, 2020), many of our participants who primarily read on screens miss the era of physical papers.
@@ 4.3.2. Faster publication through arXiv
The digital materiality of the research PDF affords an alternative publication process that came up in several of our interviews (P4,P5,P8): the preprint server arXiv. P5 says arXiv is a major source of anxiety for his students: “So my students…had the unpleasant experience of, you know, finishing a paper when they’re just about to submit and they saw a paper on arXiv that did almost the same thing. Like, oh my, months of work just went down the drain.” This emotion, the sense of loss when a project becomes unpublishable because someone else got there first, echoes Su and Crandall’s observation of “selective amnesia” (Su and Crandall, 2021), except in addition to old papers quickly becoming obsolete because something else is better, current papers may become obsolete because they are no longer first.
@@
ArXiv changes the writing process as well. As P9 explains, “I find that the amount of time and effort put into each paper has gone downhill. For my papers, I would put a huge amount of time, especially in the intro…And recently I have been doing less and less of that and basically because the students are saying, ah you know, the new kids they don’t even, they just skip the intro, they just go directly to the method. So I feel like, oh my god, nobody’s even gonna read my beautiful prose!…But also I think because the field progresses faster, papers become obsolete much quicker. So it might be reasonable not to spend so much effort on a single paper if you know that in a year it will be obsolete.” In other words, author writes differently in order to better fit the faster publication system, spending less time on introductions in order to spend less effort on a paper which might quickly become obsolete.
To summarize, the technologies underlying the publishing process in computer vision have changed rapidly over the past three decades. Today, proceedings are published online, and most researchers read research PDFs on a screen, rather than research papers in print. The loss of physical research papers affects readers’ attachment to those documents, as virtual papers cannot hold significance as domestic objects in the same way. Meanwhile, authors have taken advantage of this fact and increased the resolution of their figures to bend conference page limits, and started publishing on arXiv to quickly attract attention to their new results. Again, we see the same themes: a pattern in computer vision was preceded by a similar shift in computer graphics, which has contributed to commodification. In this case, the conference publishing system has shifted from a means of scholarly communication to a means of evaluating research. Acceptance to a computer vision conference serves as a marker, not just of peer-reviewed technical correctness, but of sufficient novelty and significance to warrant high scores from reviewers, a “stamp of approval” for employment or promotion. In other words, improvements in writing tools have shifted the burden of evaluating new research work from the authors to the peer reviewers, placing significant labor burdens on the peer review process.

### v3

@@ 4.1.2. Beyond the paper: arXiv, social media and videos
Why would many authors begin to engage in these teaser figure and title practices over the course of the 2010s? One explanation is a change in medium through which research reaches its audience, from conference and journal proceedings onto social media and preprint websites. P4 describes how his students have strategies to get their papers noticed on the preprint site arXiv: “if they take your submission on Thursday and it goes up on the arXiv most recent publications list, it’s up for all of Friday, Saturday, and Sunday…so it gets more days of exposure…It improves your odds that somebody will notice your thing.” P8 describes the pressure on her students to promote their work: “social media, like promoting one’s research has become such a big thing and I think students…are realizing, ‘Oh it’s not enough for me to, y’know, come up with a paper, post it on arXiv, get it accepted. I also need to tweet about it.’ It’s frankly quite exhausting…[before] the only way you promote your paper is it shows up at the conference and hopefully, y’know, some famous computer vision researchers will come up and look.” She talks about how her advisor would bring his friends over to her poster, and that was all the advertising she needed to get noticed, but that strategy no longer works as well in the modern crowded field of computer vision.
@@
That attention translates into citations, as well as other benefits. P11, who works in an industry lab that hires recent PhDs, explains why his group publishes at all: “exposure is actually really important. If you want to attract really top level talent, then having zero published papers is really going to work against you, right? Particularly if you’re looking at people who are in positions that are for doing some sort of active innovation.” For this group, the publication is primarily a way to gather attention, they derive value from the attention paid to their work through hiring. Second, attention helps researchers to identify the value of their own work. P7 talks at length about his desire for impact. He switched fields from computational complexity to computer vision because there were more active researchers, and he could not tell if his work was having an impact in a small field. P9 tells the story of how attention from a senior scholar improved his relationship with his advisor: “that email from [scholar], that’s the beginning of my great relationship with my advisor was that one little email, because within five minutes [advisor] was up in the office, ‘Oh, you got an email from [scholar]? Very good. Very good. How about a coffee?’ ” Notice that the career benefits resulted not from the attention itself, but from the record of that attention in the form of an email from a senior scholar.
While the relationship between attention and prestige is not new, the unprecedented growth of computer vision over the past decade places pressure on the attention of senior researchers. There have been ongoing discussions in the computer vision community on how to react to this growth. Some of the proposed ideas are recorded in the passed motions of the annual meeting of the IEEE pattern analysis and machine intelligence technical committee (PAMI-TC), the governing body of the computer vision research community. In 2021–2022, motions changed the rules regarding double submission to reduce burden on reviewers, introduced punitive measures for missing or poor quality peer reviews from authors, and banned discussion of submitted papers on social media to avoid influencing peer reviewers (the last of which was repealed in 2023) (pam, 2024; Black, 2021). The rationale statement for the social media ban specifically highlights attention: “Groups with large followings and the resources to mount visible social media promotions received significant attention for work that is under review. Reviewers are exposed to this work and the attention it receives can bias their judgment—if so many people on social media are excited, mustn’t it be good?” (Black, 2021). In addition to highlighting the valorizing role of attention, this statement showcases the view of peer review as an unbiased process for evaluating the quality of research.
@@ 4.2. Measuring the Contribution
But all of this was not the case a few decades ago. The vast majority of computer vision papers from the 1980s and 1990s rely on mathematical arguments based on, for example, pinhole camera geometry and do not contain any quantitative results. Empirical evaluation, if included at all, was primarily qualitative, in the form of figures showing sample results. As P8 explains, she had a combination of quantitative and qualitative evaluation in her paper from 2003, which was unusual: “quantitative evaluation, you know, back in 2003 was still kind of in its infancy…I’m not sure that this [2003] paper has basically any comparison to competing methods which probably would be required today.” P3 explains that he was primarily concerned in 1990 with showing test examples to demonstrate his algorithm’s effective handling of edge cases. P9 explains that in 1999, showing example output of his system was sufficient: “instead of [Amazon] Mechanical Turk you just have the reviewers just eyeball the images.” In the satirical 2010 “Paper Gestalt” (Von Bearnensquash, 2010) paper, which attempts to use computer vision methods to distinguish between good and bad papers, large confusing tables were identified as a key feature of bad papers, not an essential feature of good ones.
@@
The publication of Krizhevsky, Sutskever, and Hinton’s “ImageNet Classification with Deep Convolutional Neural Networks” in 2012 (Krizhevsky et al., 2012) marks a turning point for empirical evaluation. This paper is historically significant for setting off the deep learning revolution, and its design and writing served as a foundation for the thousands of deep learning-based computer vision papers that followed. The paper’s central argument is that several “new and unusual features” lead deep convolutional neural networks to significantly outperform other methods. These features include rectified linear units (ReLU), GPU-based training, and regularization techniques like data augmentation and dropout. Crucially for our story, however, this argument is made by way of a table, shown in Figure 5 (c), with the best performance in bold. Neural network papers are obligated to use empirical evaluation, as there are insufficient theoretical guarantees for these models and they are difficult to evaluate otherwise. Over the following years, many papers followed Krizhevsky et al.’s lead, showing that deep convolutional neural networks outperform existing methods on other central problems like object detection and semantic segmentation. We can see a corresponding increase in both the average number of tables per paper, as well as the fraction of papers containing at least one table in Figure 4. While the prevalence of figures has remained relatively constant, tables have become significantly more common. While only 75% of CVPR 2013 papers had a table, 95% of CVPR 2021 papers did, and the average number of tables per paper has doubled from two to four. While not all of these tables are the kind of results table we are discussing, the change is striking.
@@
More broadly, in computer vision there is a widespread assumption that research inherently involves competition between technical methods. Several of our participants described academic publication using free market metaphors. P2 explains his research output: “publication was really fast during that time because there was not a lot of competition.” P3 explained that CVPR has become inaccessible to students because “supply and demand” have raised the standards for publication. P4 described vision as “very industrial” and gave examples of techniques that students use to optimize their arXiv submissions to reach as many eyes as possible.
With these tables buried deep within the paper, it may seem like their effect on attention is minimal. However, a competitive benchmark result attracts attention on its own. Once a researcher has the lead, other papers must compare against their benchmark and cite them. P6 compares benchmarks to arcade game leaderboards: “If it’s an established benchmark… there’s somebody who has the lead right? Like you would have on like a classic video game arcade machine, right? It’s like I have to just get my initials at the top right? That’s exactly what they’re doing, and that’s frustrating.” P5 and P8 advise their students to avoid working on research which forces them into competition with large companies. P5 explains, “I tell them, don’t work on problems that, you know, a lot of people are working on right now, you can’t possibly compete with Facebook, Google, Amazon because you’re not as computing heavy.” While students could still develop innovative ideas for those problems, they likely would not beat the benchmarks that large companies have set without leveraging similar computing power and would not attract attention. Because students will not be able to compete with large companies for attention on these problems, they are not worth studying.
In summary, the development of the results table shows the transition of computer vision from an applied mathematical discipline to a quantitative-empirical one. As vision started relying on empirical evaluation, it adopted a style of results table which was used for runtime benchmarking in computer graphics, and as evaluation benchmarks became established for vision problems, these tables grew in size and importance. Today, competition on benchmarks is the organizing principle of the discipline; new methods must demonstrate that they are empirically more effective than existing methods to be accepted, and the results table is an essential part of the paper. This element showcases several of the same themes as teaser images: influence from computer graphics, and commodification, in the form of measurable improvements over prior work. It also echoes patterns observed in HCI more broadly regarding quantification: once a phenomenon, in this case the quality of a method, has been measured, that measurement creates and constrains possibilities for action (Pine and Liboiron, 2015).
@@ 4.3.1. The PDF and digital proceedings
The final element we explore is the Portable Document Format (PDF) itself, and the way that authors engage with it as a medium. PDF was a derivative of Adobe’s already-successful Postscript language for specifying documents, and a proprietary file format until 2008. In an article from 1998, Kasdorf compares the PDF format to SGML (Standard Generalized Markup Language), and argues that both formats should be used on the web: markup languages for screen-based content and page-based languages for printed content (Kasdorf, 1998).
@@
One such choice is the high resolution figure, which exploits the scalable nature of vector graphics and the zoom feature of PDF viewers to put more data in a PDF without exceeding the page limit. Several examples of this trend from CVPR 2021 are shown in Figure 6. While these figures are not overly complex or poorly designed, they are too small to be readable on a 300-dpi printed page.
@@
Several of our participants mentioned these high resolution figures. P6 and P10 are frustrated that they cannot print some PDFs because the figures are too small. P10 ascribes these figures to the increasing complexity of research: “the tendency of grouping more small figures into a kind of collection… I wonder, is this because of the limited page count that people try to avoid the whitespace around the figures and then they put everything in one big figure because then you can save on the whitespace? Or is it simply because really…the techniques that we’re describing nowadays are so, so much more complicated? So you need them to…put the architecture of the whole thing in detail there. Otherwise nobody understands it.” P7, on the other hand, ascribes it to digital submissions: “In ’94, you turned in your paper by actually having a piece of large paper that was probably 24” wide by 30” or something…you had columns of text from LaTeX and then your figures had to be glued in place in what’s called ‘screen,’ right? So you actually have to make sure that they didn’t copy badly.” When papers were submitted by mail, the medium of the “camera-ready” draft was a significant limitation on the kinds of visual content that could be included, but now the only limitation is on the number of pages, so authors use smaller figures to include more content. The small size of these figures, especially in comparison with the large front page teaser image, creates a visual hierarchy, highlighting the differences in purpose between these figures. The teaser attracts attention, the dense figures provide details for those who zoom in, and the impression of details for those who do not.
@@
The digital nature of contemporary research papers was a sensitive topic in our interviews. Half of our participants (P1, P2, P3, P4, P5, and P8) expressed attachment to and nostalgia for their printed papers, and were saddened when asked when they stopped reading on paper. For P1, the time spent at the library made the papers physically significant: “[the papers] were all me standing at a Xerox machine with a journal…I didn’t find it as annoying. I found that it gave me something like a kind of a physical connection to the paper. This was sort of like, it’s mine. I’ve watched it go by, you know.” P2 remembers that the SIGGRAPH proceedings “were like flipped to the point where like it’s the threads were like bare.” He’s kept all of the papers he printed during his master’s degree and revisits them from time to time: “I still go back, you know, look at like oh this crap that I thought about these things, right?” Finding and photocopying papers made them precious; he thinks there is “a kind of a correlation between that sense of, like scarcity of it, like how precious the paper was, because it’s so hard to find versus the amount of care you give it.” P4 fondly recalls his advisor taking papers out of a filing cabinet and photocopying them for him: “the amazing thing about your advisor is you have this filing cabinet of stuff that’s already pre-copied, you know, every paper he thought was interesting.”
@@
These sorts of relationships to printed papers are interesting from a design perspective. They echo the findings of Odom et al. that we preserve designed objects which have functional, symbolic and material-aesthetic value (Odom et al., 2009). When a researcher prints out and writes on a paper, the paper gains symbolic significance, a “physical connection” which endures over time. That makes printed pages less disposable than digital ones; there is a more tangible opportunity cost to creating them. From this perspective, digitization is an essential component of commodification, as it separates the research from its paper and thus its role as symbolic object. While reading digitally is significantly more environmentally sustainable than printing, shipping, and photocopying paper (Aydemir and Özsoy, 2020), many of our participants who primarily read on screens miss the era of physical papers.
@@ 4.3.2. Faster publication through arXiv
The digital materiality of the PDF affords an alternative publication process that came up in several of our interviews (P4, P5, P8): the preprint server arXiv. P5 says arXiv is a major source of anxiety for his students: “So my students…had the unpleasant experience of, you know, finishing a paper when they’re just about to submit and they saw a paper on arXiv that did almost the same thing. Like, oh my, months of work just went down the drain.” This emotion, the sense of loss when a project becomes unpublishable because someone else got there first, echoes Su and Crandall’s observation of “selective amnesia” (Su and Crandall, 2021), except in addition to old papers quickly becoming obsolete because something else is better, current papers may become obsolete because they are no longer first.
@@
ArXiv changes the writing process as well. As P9 explains, “I find that the amount of time and effort put into each paper has gone downhill. For my papers, I would put a huge amount of time, especially in the intro…And recently I have been doing less and less of that and basically because the students are saying, ah you know, the new kids they don’t even, they just skip the intro, they just go directly to the method. So I feel like, oh my god, nobody’s even gonna read my beautiful prose!…But also I think because the field progresses faster, papers become obsolete much quicker. So it might be reasonable not to spend so much effort on a single paper if you know that in a year it will be obsolete.” In other words, the author now writes differently in order to better fit the faster publication system, spending less time on introductions in order to spend less effort on a paper which might quickly become obsolete.
To summarize, the technologies underlying the publishing process in computer vision have changed rapidly over the past three decades. Today, proceedings are published online, and most researchers read PDFs on a screen, rather than research papers in print. The loss of physical research papers affects readers’ attachment to those documents, as virtual papers cannot hold significance as domestic objects in the same way. Meanwhile, authors have taken advantage of this fact and increased the resolution of their figures to bend conference page limits, and started publishing on arXiv to quickly attract attention to their new results. Again, we see the same themes: a pattern in computer vision was preceded by a similar shift in computer graphics, which has contributed to commodification. In this case, the conference publishing system has shifted from a means of scholarly communication to a means of evaluating research. Acceptance to a computer vision conference serves as a marker, not just of peer-reviewed technical correctness, but of sufficient novelty and significance to warrant high scores from reviewers, a “stamp of approval” for employment or promotion. In other words, improvements in writing tools have shifted the burden of evaluating new research work from the authors to the peer reviewers, placing significant labor burdens on the peer review process.